OpenStudy (anonymous):

Simplify the following expression. log(25x^3) + 3 log(1/x).

OpenStudy (anonymous):

how can u rewrite 3log(1/x)?

OpenStudy (anonymous):

log(25x^3) + 3 log(1/x) = log25 + log(x^3) - 3 log x = log25 +3 log x - 3 logx = log25